What will be the change of temperature for 300 gr of iron (Fe) which have gained 2000 cal? (cFe= 0.11 cal/groC)


1
Expert's answer
2021-02-06T10:05:15-0500
"Q=mc\Delta T,""\Delta T=\dfrac{Q}{mc}=\dfrac{2000\ cal}{300\ g\cdot 0.11\ \dfrac{cal}{g \cdot^{\circ}C}}=60.6\ ^{\circ}C."